OpenStudy (anonymous):

I need help with one more problem. x+3y=6 4x-2y=-32 (-6,4)?

OpenStudy (anonymous):

Yes.

OpenStudy (adunb8):

are you looking for x?

OpenStudy (anonymous):

She already found them.

OpenStudy (adunb8):

oh shes asking if those are the answers?

OpenStudy (anonymous):

(x,y). She better get 100% on her test. =]

OpenStudy (adunb8):

haha isee

OpenStudy (anonymous):

No i need help on how to get that answer.

OpenStudy (anonymous):

How'd you get the answer then?

OpenStudy (anonymous):

\[-4[ x + 3y = 6 ] =\frac{ -4x - 12y = -24 }{ 4x -2y =-32 }\] \[-14y = -56 \] \[y = 4 \] Then plug y =4 in one of the original equations and you'll find x.

OpenStudy (anonymous):

My teacher, but idk how she got it. The way she explained it didn't make sense.

OpenStudy (anonymous):

Multiply one equation so that one of it's terms cancels with the second equation.

OpenStudy (anonymous):

@xmarioz25 stop doing problems for people http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Oh, my bad. Just joined a few hours ago.

OpenStudy (anonymous):

He didn't do it for me. He just showed me how to do the first part and got me started.

OpenStudy (anonymous):

Thanks Mario (: I got it from here.

OpenStudy (anonymous):

This is like the fourth question you've asked and you still don't seem to have any idea how to solve them. Either use the site to learn or don't use it.

OpenStudy (anonymous):

Maybe it's because I'm dumb, why don't you just back off. I don't need your judgment.

OpenStudy (anonymous):

If I wanted an "answer" and not to "learn" I would have just asked yahoo, but I wanted to learn how to do the problems so I came here.

OpenStudy (anonymous):

It's not 'my' judgment you fake profile pic, it's the rules of use for the site.